# FreeReg: Image-to-Point Cloud Registration Leveraging Pretrained Diffusion Models and Monocular Depth Estimators

### v1

@@ Abstract
Matching cross-modality features between images and point clouds is a fundamental problem for image-to-point cloud registration. However, due to the modality difference between images and points, it is difficult to learn robust and discriminative cross-modality features by existing metric learning methods for feature matching. Instead of applying metric learning on cross-modality data, we propose to unify the modality between images and point clouds by pretrained large-scale models first, and then establish robust correspondence within the same modality. We show that the intermediate features, called diffusion features, extracted by depth-to-image diffusion models are semantically consistent between images and point clouds, which enables the building of coarse but robust cross-modality correspondences. We further extract geometric features on depth maps produced by the monocular depth estimator. By matching such geometric features, we significantly improve the accuracy of the coarse correspondences produced by diffusion features. Extensive experiments demonstrate that without any task-specific training, direct utilization of both features produces accurate image-to-point cloud registration. On three public indoor and outdoor benchmarks, the proposed method averagely achieves a $20.6\%$ improvement in Inlier Ratio, a 3.0×3.0$\times$3.0 × higher Inlier Number, and a $48.6\%$ improvement in Registration Recall than existing state-of-the-arts. The codes and additional results are available at https://whu-usi3dv.github.io/FreeReg/.
@@ 2 Related work
Image-to-point cloud registration. In order to establish correspondences between images and point clouds for pose recovery, most existing methods (Li et al., 2015; Xing et al., 2018; Feng et al., 2019; Lai et al., 2021; Wang et al., 2021; Pham et al., 2020; Liu et al., 2020; Jiang & Saripalli, 2022; Li et al., 2023) rely on metric learning to align local features of images and point clouds (Feng et al., 2019; Pham et al., 2020; Lai et al., 2021; Jiang & Saripalli, 2022), or depth maps (Liu et al., 2020; Wang et al., 2021; Li et al., 2023). However, these methods often require cross-modal registration training data (Pham et al., 2020; Wang et al., 2021; Jiang & Saripalli, 2022; Li et al., 2023) and show limited generalization ability (Pham et al., 2020; Wang et al., 2021; Li et al., 2023) due to the difficulty in the cross-modality metric learning. In contrast, FreeReg does not require task-specific training and finetuning and exhibits strong generalization ability to both indoor and outdoor scenes.
Some other methods directly solve image-to-point cloud registration as an optimization problem (David et al., 2004; Campbell et al., 2019), which regresses poses by progressively aligning keypoints (Li & Lee, 2021; Ren et al., 2022; Campbell et al., 2019), pole structures (Wang et al., 2022b), or semantic boundaries (Liao et al., 2023) of RGB images and depth maps. However, these methods heavily rely on an accurate initial pose (Wang et al., 2021; Liao et al., 2023) to escape from local minima in optimizations. Thus, these methods are mostly constrained to specific application scenarios (Li & Lee, 2021; Ren et al., 2022; Arar et al., 2020). FreeReg does not require such a strictly accurate initialization because FreeReg matches features to build correspondences to handle large pose changes.
@@
Diffusion feature for matching. Some recent works utilize diffusion features for representation learning (Kwon et al., 2022) and semantic matching (Luo et al., 2023; Tang et al., 2023; Hedlin et al., 2023; Zhang et al., 2023) among RGB images capturing objects across instances and categories. In comparison, our method shows the effectiveness of diffusion features in learning cross-modality features for image-to-point cloud registration.
@@ 3.2 Diffusion Features on cross-modality data
RGB diffusion feature. As shown in Fig. 4(a), we perform the forward process of SD (Rombach et al., 2022) to add noise to the input RGB image, which results in a noisy image on the same predefined step $\hat{t}$. The noisy image is fed to the UNet of the SD and the intermediate feature maps of the UNet decoder are used as the diffusion feature for the input RGB image.
Depth diffusion feature. Given the depth maps, we first densify them using traditional erosion and dilation operations (Ku et al., 2018). As shown in Fig. 4 (b), we propose to feed the depth map to a CN (Zhang & Agrawala, 2023) as a condition to guide the reverse process of SD. With such a condition, SD gradually denoise a pure Gaussian noise until a predefined step $\hat{t}$ and then we use the feature maps in the SD UNet decoder as the depth diffusion features. An alternative way is to directly treat the depth map as an RGB image for diffusion feature extraction, which however leads to poor performance as shown in the supplementary material.
@@ 4.1 Experimental protocol
Metrics. Following (Choy et al., 2019; Wang et al., 2023b; a), we adopt four evaluation metrics: (1) Feature Matching Recall (FMR) is the fraction of I2P pairs with more than $5\%$ correct estimated correspondences. A correspondence is regarded as correctly matched if its ground truth 3D distance is smaller than $\tau_{c}$. $\tau_{c}$ is set to 0.3m for 3DMatch/ScanNet and 3m for Kitti-DC. (2) Inlier Ratio (IR) is the average correct correspondence proportions among all I2P pairs. (3) Inlier Number (IN) is the average number of correct correspondences on each I2P pair. and (4) Registration Recall (RR) is the percentage of correctly-aligned I2P pairs with rotation and translation errors less than $\tau_{R}$ and $\tau_{t}$ respectively. ($\tau_{R}$, $\tau_{t}$) is set to ($20^{\circ}$, 0.5m) for 3DMatch/ScanNet and ($10^{\circ}$, 3m) for Kitti-DC. We provide additional results under different threshold conditions in the supplementary material.
Baselines. We compare FreeReg with fully supervised registration baselines. The image registration method SuperGlue (SG) (Sarlin et al., 2020) is modified to match RGB images and point clouds. LCD (Pham et al., 2020) learns to construct I2P cross-modality descriptors utilizing metric learning. DeepI2P (Li & Lee, 2021) resolve I2P registration by optimizing an accurate initial pose. We implement a cross-modality feature extraction method I2P-Matr following a concurrent work 2D3D-Matr (Li et al., 2023), where the official codes are not released yet. Meanwhile, we compare FreeReg with P2-Net (Wang et al., 2021) and 2D3D-Matr (Li et al., 2023) under their experimental protocol (Li et al., 2023) in the supplementary material, where FreeReg also achieves the best registration performance. We also adopt a combined baseline which first utilizes ControlNet (Zhang & Agrawala, 2023) (CN+SG) to generate an RGB image from the target point cloud and then conducts SuperGlue (Sarlin et al., 2020) to match the input and the generated image. For our method, we report the results using only the diffusion feature (FreeReg-D), only the geometric feature (FreeReg-G), and the fused feature (FreeReg) for matching. More implementation and experimental details are provided in the supplementary material.
@@ 4.2 Results on three benchmarks
4.3 Ablation studies
We conducted comprehensive ablation experiments on FreeReg. More ablation studies on diffusion feature extraction and I2P transformation estimation are provided in the supplementary material.
@@ 4.3.1 Ablating diffusion feature extraction
Diffusion step selection. In Table 3, we aim to determine the diffusion step $\hat{t}$. The experimental results demonstrate that the Diffusion Features from $\hat{t}=150$ achieve the best registration performance on BFO. Results on 3DMatch confirm its effectiveness.
@@ 4.3.2 Ablating feature fusion weight
We ablate the fusion weight $w$ to fuse diffusion and geometric features in Table. 4 based on the baseline model FreeReg. It can be seen that FreeReg achieves the best registration performance when $w$ is set to 0.5. Moreover, we find that relying more on diffusion features, i.e., $w=0.6$ achieves a much similar result to the default FreeReg. While relying more on geometric features, i.e., $w=0.4$ causes a sharp performance drop of a $8.7\%$ lower IR and a $2.5\%$ lower RR. This demonstrates the robustness of the proposed diffusion features.
@@ 4.4 Limitations
The main limitation is that FreeReg requires about 11s and 13G GPU memory to match a single I2P pair on a 4090 GPU. The reason is that we need to run multiple backward process steps of ControlNet to denoise the pure noises to reach a specific step $\hat{t}$ for feature extraction. Meanwhile, though we show the superior performance of using diffusion features for I2P registration, we manually select layers and denoising steps in the diffusion feature extraction, which could be improved by future works to automatically select good features.

### v2

@@ Abstract
Matching cross-modality features between images and point clouds is a fundamental problem for image-to-point cloud registration. However, due to the modality difference between images and points, it is difficult to learn robust and discriminative cross-modality features by existing metric learning methods for feature matching. Instead of applying metric learning on cross-modality data, we propose to unify the modality between images and point clouds by pretrained large-scale models first, and then establish robust correspondence within the same modality. We show that the intermediate features, called diffusion features, extracted by depth-to-image diffusion models are semantically consistent between images and point clouds, which enables the building of coarse but robust cross-modality correspondences. We further extract geometric features on depth maps produced by the monocular depth estimator. By matching such geometric features, we significantly improve the accuracy of the coarse correspondences produced by diffusion features. Extensive experiments demonstrate that without any training on the I2P registration task, direct utilization of both features produces accurate image-to-point cloud registration. On three public indoor and outdoor benchmarks, the proposed method averagely achieves a $20.6\%$ improvement in Inlier Ratio, a 3.0×3.0$\times$3.0 × higher Inlier Number, and a $48.6\%$ improvement in Registration Recall than existing state-of-the-arts. The code and additional results are available at https://whu-usi3dv.github.io/FreeReg/.
@@ 2 Related work
Image-to-point cloud registration. In order to establish correspondences between images and point clouds for pose recovery, most existing methods (Li et al., 2015; Xing et al., 2018; Feng et al., 2019; Lai et al., 2021; Wang et al., 2021; Pham et al., 2020; Liu et al., 2020; Jiang & Saripalli, 2022; Li et al., 2023) rely on metric learning to align local features of images and point clouds (Feng et al., 2019; Pham et al., 2020; Lai et al., 2021; Jiang & Saripalli, 2022), or depth maps (Liu et al., 2020; Wang et al., 2021; Li et al., 2023). However, these methods often require cross-modal registration training data (Pham et al., 2020; Wang et al., 2021; Jiang & Saripalli, 2022; Li et al., 2023; Kim et al., 2023) and show limited generalization ability (Pham et al., 2020; Wang et al., 2021; Li et al., 2023; Ren et al., 2022; Yao et al., 2023) due to the difficulty in the cross-modality metric learning. In contrast, FreeReg does not require any training and fine-tuning on the I2P registration task and exhibits strong generalization ability to both indoor and outdoor scenes.
Some other methods directly solve image-to-point cloud registration as an optimization problem (David et al., 2004; Campbell et al., 2019; Arar et al., 2020; Wang et al., 2023a; Zhou et al., 2023), which regresses poses by progressively aligning keypoints (Li & Lee, 2021; Ren et al., 2022; Campbell et al., 2019), pole structures (Wang et al., 2022b), semantic boundaries (Liao et al., 2023), or cost volumes (Wang et al., 2023a) of RGB images and depth maps. However, these methods heavily rely on an accurate initial pose (Wang et al., 2021; Liao et al., 2023) to escape from local minima in optimizations. FreeReg does not require such a strictly accurate initialization because FreeReg matches features to build correspondences to handle large pose changes.
@@
Diffusion feature for matching. Some recent works utilize diffusion features for representation learning (Kwon et al., 2022) and semantic matching (Luo et al., 2023a; Tang et al., 2023; Hedlin et al., 2023; Zhang et al., 2023) among RGB images capturing objects across instances and categories. In comparison, our method shows the effectiveness of diffusion features in learning cross-modality features for image-to-point cloud registration.
@@ 3.2 Diffusion Features on cross-modality data
RGB diffusion feature. As shown in Fig. 4(a), we perform the forward process of SD (Rombach et al., 2022) to add noise to the input RGB image, which results in a noisy image on a predefined step $\hat{t}$. The noisy image is fed to the UNet of the SD and the intermediate feature maps of the UNet decoder are used as the diffusion feature for the input RGB image.
Depth diffusion feature. Given the depth maps, we first densify them using traditional erosion and dilation operations (Ku et al., 2018). As shown in Fig. 4 (b), we propose to feed the depth map to a CN (Zhang & Agrawala, 2023) as a condition to guide the reverse process of SD. With such a condition, SD gradually denoise a pure Gaussian noise until the same predefined step $\hat{t}$ and then we use the feature maps in the SD UNet decoder as the depth diffusion features. An alternative way is to directly treat the depth map as an RGB image for diffusion feature extraction, which however leads to poor performance as shown in the supplementary material.
@@ 4.1 Experimental protocol
Metrics. Following (Choy et al., 2019; Wang et al., 2023c; b), we adopt four evaluation metrics: (1) Feature Matching Recall (FMR) is the fraction of I2P pairs with more than $5\%$ correct estimated correspondences. A correspondence is regarded as correctly matched if its ground truth 3D distance is smaller than $\tau_{c}$. $\tau_{c}$ is set to 0.3m for 3DMatch/ScanNet and 3m for Kitti-DC. (2) Inlier Ratio (IR) is the average correct correspondence proportions among all I2P pairs. (3) Inlier Number (IN) is the average number of correct correspondences on each I2P pair. and (4) Registration Recall (RR) is the percentage of correctly-aligned I2P pairs with rotation and translation errors less than $\tau_{R}$ and $\tau_{t}$ respectively. ($\tau_{R}$, $\tau_{t}$) is set to ($20^{\circ}$, 0.5m) for 3DMatch/ScanNet and ($10^{\circ}$, 3m) for Kitti-DC. We provide additional results under different threshold conditions in the supplementary material.
Baselines. We compare FreeReg with fully supervised registration baselines. The image registration method SuperGlue (SG) (Sarlin et al., 2020) is modified to match RGB images and point clouds. LCD (Pham et al., 2020) learns to construct I2P cross-modality descriptors utilizing metric learning. DeepI2P (Li & Lee, 2021) resolve I2P registration by optimizing an accurate initial pose. We implement a cross-modality feature extraction method I2P-Matr following a concurrent work 2D3D-Matr (Li et al., 2023), where the official codes are not released yet. Meanwhile, we compare FreeReg with P2-Net (Wang et al., 2021) and 2D3D-Matr (Li et al., 2023) under their experimental protocol (Li et al., 2023) in the supplementary material, where FreeReg also achieves the best registration performance. We also adopt a baseline as mentioned in Fig. 2 (I) that first utilizes ControlNet (Zhang & Agrawala, 2023) to generate an RGB image from the target point cloud and then conducts SuperGlue (Sarlin et al., 2020) to match the input and the generated image (CN+SG). For our method, we report the results using only the diffusion feature (FreeReg-D, i.e. $w=1$), only the geometric feature (FreeReg-G, i.e. $w=0$), and the fused feature (FreeReg, i.e. $w=0.5$ by default) for matching. More implementation details and analysis are provided in the supplementary material.
@@ 4.2 Results on three benchmarks
4.3 More analysis
We conduct comprehensive ablation experiments on FreeReg in the following. More analysis about how to fine-tune FreeReg for performance gains, runtime analysis and acceleration strategies of FreeReg, FreeReg performance under different thresholds, more qualitative comparisons and ablation studies are provided in the supplementary material.
@@ 4.3.1 Ablating diffusion feature extraction
Diffusion step selection. In Table LABEL:tab:step, we aim to determine the diffusion step $\hat{t}$. The experimental results demonstrate that the Diffusion Features from $\hat{t}=150$ achieve the best registration performance on BFO. Results on 3DMatch confirm its effectiveness.
@@ 4.3.2 Ablating feature fusion weight
We ablate the fusion weight $w$ to fuse diffusion and geometric features in Table. LABEL:tab:fusew based on the baseline model FreeReg. It can be seen that FreeReg achieves the best registration performance when $w$ is set to 0.5. Moreover, we find that relying more on diffusion features, i.e., $w=0.6$ achieves a much similar result to the default FreeReg. While relying more on geometric features, i.e., $w=0.4$ causes a sharp performance drop of a $8.7\%$ lower IR and a $2.5\%$ lower RR. This demonstrates the robustness of the proposed diffusion features.
@@ 4.4 Limitations
The main limitation is that FreeReg requires about 9.3s and 12.7G GPU memory to match a single I2P pair on a 4090 GPU, yielding much higher RR but longer time usage than baselines LCD (0.6s, 3.5G, I2P-Matr (1.7s, 2.7G), and CN+SG (6.4s, 11.6G). The reason is that we need to run multiple backward process steps of ControlNet to denoise the pure noises to reach a specific step $\hat{t}$ for feature extraction. In the supplementary material, we show how to accelerate FreeReg by $\sim 50\%$ with only a $\sim 1.4\%$ RR drop. Meanwhile, though we show the superior performance of using diffusion features for I2P registration, we manually select layers and denoising steps in the diffusion feature extraction, which could be improved by future works to automatically select good features.